OpenStudy (anonymous):

The length of a rectangular box is 2 units less than its width. If the area of the box is 15 units square. Find the perimeter of the box. 15=w(2-w)

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Im trying to find the perimeter of that rectangular box. But I need to find out the width first using that formula I gave.

OpenStudy (anonymous):

the answer is 16 Explanation These are the formulas for the box. A= L x W P = 2( L + W) We know that the area is 15 15= L x W Also, L= W-2 Now we can use both formulas (W)(W-2) = 15 W² - 2W= 15 W²- 2W-15= 0 (W-5) (W+3) = 0 W= 5 , W= -3 We can find the Length, L = W-2 = 5- 2= 3 L= 3 Finally the perimeter: P = 2( 3+ 5) P = 2(8) P= 16